OpenStudy (anonymous):

Which of the following equations represents a line that is perpendicular to the line that passes through the points below? (-7,-4) (-5,2)

OpenStudy (tkhunny):

You'll need the slope of the line containing those points. Do it!

OpenStudy (anonymous):

A. y=-1/3x+2 B. y=1/3x+2 C. y=-3x+2 D. y=3x+2

OpenStudy (tkhunny):

Oh, well in that case, you'll need the slope of the line containing those points. Do it!